OpenStudy (anonymous):

Suppose p=x/n, where is a sample proportion, x is the number of people in the sample who share a particular characteristic, and n is the total number of people in the sample. If p=0.6 and x=24 , what is n ?

OpenStudy (amistre64):

hmm, what would you propose is a good approach to a solution?

OpenStudy (amistre64):

if p = x/n, how do we solve for n?

OpenStudy (anonymous):

i divided 0.6 by 24 and got 0.025 but i dont know what else to do

OpenStudy (amistre64):

well, thats fine, but what you have right now is p/x = 1/n youve got a reciprocal of n. lets do this p = x/n , multiply both sides by n np = x, now we can divide both sides by p n = x/p now this solves for n, instead of 1/n

OpenStudy (amistre64):

what is 24/.6 ?

OpenStudy (anonymous):

its 40

OpenStudy (amistre64):

thats better notice that: 1/40 = .025 , what you got at first. 40/1 = 40 whichis what we need.

OpenStudy (anonymous):

you made it seem so easy, the online class make everything confusing